OpenStudy (anonymous):

Okay, I am studying for a test for college, I am stuck on some of these questions and not only need an answer, but a little bit of an explanation to understand so I can take this test, please. Change each of the following measurements into an equivalent value in standard exponential notation, having the proper number of significant digits. A. 0.00322 cm B. 9901 m C. 176.0mL Please explain as well as you can.

OpenStudy (abb0t):

Well, anything that you are moving to the \(right\), you know you should know that your exponential power is going to be \(\sf \color{red}{negative}\). For example, if you have 10\(^{-3}\) = \(\sf \frac{1}{10^3}\) = 0.001

OpenStudy (abb0t):

So, for a.) you would move it to after the 3, making sure to keep significant figures. So you would have: 3.22 \(\sf \times 10^{-3}\) You try b.

OpenStudy (anonymous):

So B would be \[9.901 x 10^{3}\] ?

OpenStudy (anonymous):

Also, on C. would you keep the 0 and it would be 1.760 X 10^2 or would the 0 drop? This is what confuses me the most.

OpenStudy (abb0t):

Yes. You would keep the zero because it is a significant figure.

OpenStudy (anonymous):

Okay. can you help me on a few more? And if so, do you want me to post them as a separate question or on this question?

OpenStudy (abb0t):

Here is fine. I need to get off soon though, so try to post it soon.

OpenStudy (anonymous):

How many significant numbers does the following contain? 8.5201+ 1.93 that's one of the questions, when you have to go I will just post them separate for others to help.

OpenStudy (abb0t):

You like them up together.

OpenStudy (abb0t):



OpenStudy (anonymous):

And then would I add them?

OpenStudy (abb0t):

Round the answer to the LEAST number of places in the decimal portion of any number in the problem.

OpenStudy (abb0t):

So you would have something like: \(\sf \color{red}{n.pq}\) were each letter represents a number.

OpenStudy (anonymous):

Okay. so my answer when adding was 10.4501 would I round it to 10.45 or 10.5?

OpenStudy (abb0t):

45

OpenStudy (anonymous):

Okay. Thank you. Would it be the same for How many significant numbers does the difference 218.7201-218.63 contain?